OpenStudy (vshiroky):

While walking to school, Aaron kept finding money. When he got to school he counted it and found that he had 7/12 of 15 dollars. in simplest terms, how much money did he have?

OpenStudy (vshiroky):

@ganeshie8 can you help with this last one?

OpenStudy (uglyturtle):

He had 8.75 dollars. (8 dollars and 75 cents)

OpenStudy (vshiroky):

I have to be able to explain it.

OpenStudy (uglyturtle):

7/12 = 0.58333333333 Of = Times 16 = 15 0.58333333333 x 15 = 8.75 8.75 dollars Or 8 dollars and 75 cents.

OpenStudy (vshiroky):

I don't know why, but I am still not getting it lol

ganeshie8 (ganeshie8):

okay, so u wanto simplifu : "7/12 of 15 dollars"

ganeshie8 (ganeshie8):

right ?

OpenStudy (vshiroky):

I guess so..

OpenStudy (vshiroky):

You would think after what I have done this would be easy lol

ganeshie8 (ganeshie8):

let me ask u a quick question, whats half of 15 dollars ?

OpenStudy (vshiroky):

7.50

ganeshie8 (ganeshie8):

that means, 1/2 of 15 dollars is 7.5 ?

OpenStudy (vshiroky):

Right

ganeshie8 (ganeshie8):

how did u get that ?

OpenStudy (vshiroky):

Lol I just knew it off the top of my head...

ganeshie8 (ganeshie8):

c'mon, give it a try... u wud done something top of ur head... :)

OpenStudy (vshiroky):

15/1*1/2?

ganeshie8 (ganeshie8):

to get 1/2 of 15 dollars, we simply multiply 1/2 and 15 1/2 * 15 7.5

OpenStudy (vshiroky):

so then 15/2

ganeshie8 (ganeshie8):

yes !

OpenStudy (vshiroky):

So could I do 7/12*15

OpenStudy (uglyturtle):

Yes

ganeshie8 (ganeshie8):

yup !

ganeshie8 (ganeshie8):

\(\large \frac{7}{12}\times 15\)

OpenStudy (uglyturtle):

"Doubt kills more dreams than failure ever will" -Karim Seddiki

ganeshie8 (ganeshie8):

\(\large \frac{7}{12}\times 15\) \(\large \frac{7}{4\times 3}\times 3\times 5\) \(\large \frac{7}{4}\times 5\) \(\large \frac{35}{4}\) \(\large 8\frac{3}{4}\)

OpenStudy (uglyturtle):

8.75

ganeshie8 (ganeshie8):

nice quote :)

OpenStudy (uglyturtle):

Thanks :)

OpenStudy (uglyturtle):

Goodnight.

ganeshie8 (ganeshie8):

gn have good slp :)

OpenStudy (vshiroky):

Thank you!!

ganeshie8 (ganeshie8):

np :) wonder what ur wuesiton means by " simplest terms"

ganeshie8 (ganeshie8):

which of the below is simple ? 1) \(\frac{35}{4}\) 2) \(8\frac{3}{4}\) 3) \(8 dollars, 75 cents\) 4) \(8.75\)

ganeshie8 (ganeshie8):

it could be anything :\

ganeshie8 (ganeshie8):

all express same thing, and "simple terms" is bit subjective i feel lol

ganeshie8 (ganeshie8):

getting wat i mean ha ?

ganeshie8 (ganeshie8):

if you're entering this answer in online grader, then you're doomed cuz you wont knw upfront wat the grader thinks as simple

OpenStudy (vshiroky):

Lol no it's for a 4th graders homework haha

ganeshie8 (ganeshie8):

ohhkie,, you helping somebody ha :)

OpenStudy (vshiroky):

Yeah my friends daughter

ganeshie8 (ganeshie8):

nice :) teaching kids is the most challenging task lol... have fun !!

OpenStudy (vshiroky):

Thanks haha